OpenStudy (anonymous):

What will the height be at t = 1 second Is that right? medals!!!

OpenStudy (anonymous):

h(t) = -16t^2 + 25t + 3 h(1) = -16(1) + 25(1) + 3 h(1) = -16 + 25 + 3 h(1) = 12 feet

OpenStudy (solomonzelman):

well, h(t) = -16t^2 + 25t + 3 `h(t) = -16(1)^2 + 25t + 3 ` h(1) = -16(1) + 25(1) + 3 h(1) = -16 + 25 + 3 h(1) = 12 feet I put the step you didn't write in gray, but it doesn't make a difference this time, since 1^2=1, so good job !! 12 is correct.

OpenStudy (anonymous):

oh ok I forget to put it ok thanks

OpenStudy (anonymous):

good night bye :)

OpenStudy (solomonzelman):

u2